OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)=4/x and g(x)= 4/x

OpenStudy (anonymous):

@hartnn can you please write out each step?

OpenStudy (anonymous):

f(g(x)) = f(4/x) = 4/(4/x) = 4x/4 = x

OpenStudy (anonymous):

did i do it right so far?

hartnn (hartnn):

yes! good :) now find g(f(x)) in same way :)

OpenStudy (anonymous):

g(f(x))= g(x)=4/x 4/(4/x) 4x/4 x

OpenStudy (anonymous):

i dont get it, did i do it right

hartnn (hartnn):

you got 'x' for both f(g(x)) and g(f(x)) that confirms that f and g are inverses of each other! and yes, you did it correctly :)

OpenStudy (anonymous):

i did the second part correctly too?

OpenStudy (anonymous):

cause i wrote the same exact thing?..

hartnn (hartnn):

its not same! in 2nd part , you actually did g(f(x)) you feel its same because both f and g are same functions!

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome ^_^